# Towards Social AI: A Survey on Understanding Social Interactions

### v1

@@ I Introduction
Social intelligence of machines to understand social interactions can be the foundation for developing various AI applications. Embodied agents [2, 3] such as robots require social understanding capabilities to effectively interact and collaborate with humans in the real world. For instance, a socially intelligent robot could pick up on verbal and non-verbal cues from people to know when to take its turn in a conversation or when to offer assistance for people. Virtual agents in simulations [4, 5] also require social understanding to engage in realistic interactions with human players. For example, virtual agents in a multiplayer game ideally should interpret content of chat messages from multiple players to infer their emotional state, intentions, and social relationships with other players. More broadly, capabilities for social understanding can enhance diverse types of AI agents, from chatbots [6] to recommendation systems [7], to better understand user needs and build rapport with humans. By integrating social understanding into systems, we can create more effective and socially adept AI applications across diverse domains.
@@
    Understanding multi-party dynamics: Social interactions often involve multiple participants with different perspectives. Analyzing group conversations requires aggregating individual social behaviors and modeling the interpersonal connections between people. This entails multifaceted aspects of analyzing dynamics such as tracking who is speaking to whom, who is interrupting whom, and who is being positive to whom. Based on the interpretation of such relationships, we can effectively understand social interactions among multiple people.
@@
As a result, the roadmap for ideal social AI involves building machines with capabilities across those three fronts: multimodal understanding, multi-party modeling, and belief awareness. Figure 1 shows the dynamics of social interactions including these three key capabilities. Through these capabilities, machines can be close to human social intelligence for seamlessly interpreting social interactions. While progress has been made in narrow applications, advanced social intelligence requires integrating breakthroughs across those areas.
This survey provides a comprehensive overview of machine learning techniques for understanding social interactions and promising future directions. By reviewing existing techniques and limitations, we aim to illuminate a path forward for endowing machines with deeper social capabilities to perceive, understand, and participate naturally in human social contexts.
We classify existing works related to understanding social interactions from the perspective of social cue types (i.e., verbal and non-verbal cues). This classification is associated with how to extract social meaning from each social cue, which can provide a foundation for holistically understanding social interactions. Figure 2 presents a taxonomy of existing research on social understanding, organized according to the types of social cues. The taxonomy includes three main branches: verbal cues, non-verbal cues, and multimodal cues. Under the verbal cues branch, key research areas include dialogue act analysis for recognizing intent, dialogue emotion analysis, and commonsense reasoning to interpret socially common knowledge in language. The non-verbal cues branch encompasses research on recognizing non-verbal behaviors including gestures, gaze, and facial expressions. Finally, the multimodal cues branch covers works that integrate signals across modalities to more comprehensively infer emotions, conversation dynamics, and social situations.
@@
The components listed under each category in Figure 2 are not meant to be isolated research areas. They represent key building blocks and capabilities that need to be connected and integrated to enable truly holistic social understanding. For example, dialogue act analysis and emotion recognition from facial expressions could be combined to better interpret the intent and sentiment behind an utterance. Similarly, gesture recognition and gaze analysis could be jointly leveraged with spoken language to recognize who an utterance is directed at in a multi-party conversation. Therefore, while organizing the landscape of existing research, the taxonomy also points towards the need for further integration across areas. In the course of our survey, we will touch upon this need for integration of multimodal social cues, as well as discuss how multi-party dynamics and belief modeling intersect with the social understanding research landscape.
@@
This is the first survey that addresses comprehensive overview of machine learning studies for social understanding, covering both verbal and non-verbal approaches. While there have been surveys that address specific aspects of social understanding such as gaze analysis [15], dialogue systems [16], and facial expression recognition [17], these works focus on narrow subdomains without considering the broader social context. There also have been several position papers on social AI regarding data infrastructure [18] and AI agents [19], which do not focus on the comprehensive survey of exisiting literatures. In contrast, our paper offers a holistic survey of the social understanding landscape, encompassing techniques across multiple modalities and identifying key capabilities that are crucial for advancing socially intelligent AI. By organizing the wide range of technical works around these core capabilities, we aim to provide a unified perspective of the field and illuminate promising research directions. The scope of our survey covers the problem of social understanding rather than creating agents that behave socially, as we argue that social understanding is a fundamental prerequisite for socially intelligent agents.
@@
We structure our paper as follows. In Sections II-IV, we provide an overview of recent advancements in machine learning research related to understanding verbal, non-verbal, and multimodal cues, respectively. Building on those sections, Section V covers benchmark datasets relevant to social understanding across modalities. Section VI discusses open challenges and promising future research directions based on the suggested key capabilities. Finally, we summarize the key points of this survey and offer concluding remarks in Section VII. By reviewing existing techniques and limitations, we aim to clarify the development trajectory and road ahead for social understanding. We hope this work crystallizes the landscape and spurs further interest around this research area.
@@
II Understanding Verbal Cues
II-A Dialogue Act Analysis
II-A1 Social Background
A dialogue act is “a combination of a communicative function and a semantic content” [21]. It refers to the intended action and communicative purpose of an utterance. During social interactions, individuals use utterances to achieve different purposes, creating impacts on their cognitive states as well as the surrounding environment defined as the dialogue context. For example, speaker A would ask speaker B for a help with opening the door, there are many possible utterances that can be used by speaker A and all these different utterances fulfil the same dialogue act. Dialogue act analysis is very important in social AI systems in terms of natural language understanding, intent detection, contextually aware and personalized interactions.
II-A2 Dialogue Act Classification
Early work treated dialogue act classification as a sequence labeling problem and applied artificial neural networks for dialogue act prediction. Lee and Dernoncourt [32] proposed a model based on convolutional neural networks (CNNs) and recurrent neutral networks (RNNs). Later work used hierarchical neutral networks for the task. Kumar et al. [33] built a hierarchical recurrent neural network using bidirectional LSTM and conditional random field which can model the dependency between utterances and labels. Such hierarchical structure is able to model representations at multiple levels across word, utterance as well as conversation levels. More recent work incorporates context-aware self attention mechanism to the hierarchical neural network to further enhance dialogue act understanding. Raheja and Tetreault [34] combined self-attention, hierarchical neural networks and contextual dependencies to develop a new model that leads to performance improvements. Li et al. [35] recognized the association between dialogue acts and subject matters and introduced dual-attention hierarchical recurrent neural network to capture such dependency. Chapuis et al. [36] used a hierarchical encoder based on transformer architectures to learn generic representations of dialogues in the pretraining stage. Colombo et al. [37] proposed a sequence-to-sequence model with a hierarchical encoder, guided attention mechanism and beam search method, which achieves state-of-the-art performance on several benchmarks. Moreover, work like [38] considered the interactive nature of dialogue by incorporating turn changes among speakers into dialogue act modeling. Malhotra et al. [20] also introduced speaker- and time-aware contextual learning in their counseling conversation dataset. As such, information about speakers can be captured and it provides more semantic information about the conversations. Furthermore, understanding individuals’ beliefs helps with dialogue state tracking. Mrksic et al. [39] leveraged this idea with a neural belief tracker to predict users’ goals in dialogues.
@@ II-A2 Dialogue Act Classification
II-A3 Discussion
Dialogue act analysis is important for understanding dialogue contexts and the information exchange among different parties. However, it often tends to focus on the local context of the dialogue exchange itself, without fully accounting for the broader social context in which the interaction is embedded. Social factors like power relations and cultural norms can shape dialogue acts in important ways that may be missed. Additionally, much of the research has focused on the verbal aspects of interaction, despite face-to-face conversations involving rich multimodal signals like gaze and gesture. Furthermore, most dialogue act analysis works are based on task-oriented or single-topic interactions between two participants, whereas real-world social interactions often involve multiple participants discussing various topics and goals, adding complexity. Analyzing dialogues with multiple participants further complicates the task as these often lack clear turn-taking and may involve overlapping speech, making it difficult to accurately identify dialogue acts.
@@ II-A3 Discussion
II-B Dialogue Emotion Analysis
II-B1 Social Background
@@
II-B2 Dialogue Emotion Recognition
Based on Poria et al. [47], the essential information required to accurately detect the emotion in an utterance from a dialogue can be categorized into three types: 1) the utterance and its surrounding context; 2) the speaker’s state; and 3) emotions in preceding utterances. Revolving around these three identified categories, researchers have come up with different methods to identify emotions embedded in conversations. Poria et al. applied bidirectional LSTM and attention mechanism to model context from surrounding utterances and Hazarika et al. introduced conversational memory network to exploit speaker memories to model the dialogue context [48]. These early works did not differentiate between various speakers. In light of this, Majumder et al. proposed DialogueRNN [49], which treats each party involved in a dialogue individually and adapts to the speaker of every utterance. Li et al. [50] framed the task under a multi-task setting by introducing an auxiliary task of speaker identification to improve emotion recognition capability in models. Besides individual speaker attributes, modeling inter-speaker relations is also very important. Yeh et al. [51] proposed Interaction-aware Attention Network (IANN) with a novel attention mechanism and Ghosal et al. [52] introduced DialogueGCN with a graph convolutional neutral network to model inter-speaker dependency. On top of inter-speaker relation modeling, researchers also use graph neutral networks to model dialogue context such as local and global context [53], as well as turn modeling [54]. Furthermore, some work incorporated common sense knowledge to enhance emotion recognition in dialogues. Ghosal et al. [55] introduced the COSMIC framework to incorporate common sense knowledge such as causal relations to better understand interactions and emotion tones. Xie et al. [56] introduced Knowledge-Interactive network which utilizes common sense knowledge and sentiment lexison to enrich semantic information. Zhu et al. [57] built a topic-augmented language model and enhanced it with common sense statements from a knowledge base grounded on the dialogue context to predict emotion labels more accurately. More recently, Lei et al. [58] framed emotion recognition in conversation to a generative plug-and-play framework using Large Language Models, which achieves state-of-the-art performance.
II-B3 Discussion
@@
II-C Common Sense Reasoning
II-C1 Social Background
@@
Common sense reasoning helps to explain the mechanisms of human mind and behavior [61, 62]. It also aids individuals explore the surroundings and understand the social context seamlessly [63]. During social interactions, common sense reasoning allows individuals to make accurate inferences about others’ beliefs based on observable evidence and social context. Ambiguity in language remains one important challenge for language understanding and common sense reasoning is one of the keys in resolving such ambiguity [64].
II-C2 Common Sense Reasoning
@@
In the past decade, neural methods have demonstrated remarkable progress in common sense reasoning tasks. In early work, embedding models such as word2vec [65] and GloVe [66] are trained to learn representations of different words and entities in the vector space. Such word representations can help models understand the relationships between different concepts, enhancing their common sense reasoning capabilities. Model architectures like memory networks [67] and Recurrent Entity Network [68] can keep track of the world state with long-term memory. In more recent years, the transformer architecture has been applied in the Generative Pre-trained Transformer (GPT) [69] and the Bidirectional Encoder Representations from Transformers (BERT) [70]. Compared to previous pretrained embedding models, they are able to learn contextual word representation with the attention mechanism. With such contextual language models, the meaning of a word can be different and hence is better captured based on the surrounding context, leading to stronger common sense reasoning capabilities in models. These pretrained models can be further fine-tuned for downstream common sense reasoning tasks to become more task-specific [71]. With increased scale in language models, recent large language models can perform common sense reasoning in zero-shot or few-shot [72] manners. Moreover, Wei et al. introduced step-by-step thinking through Chain of Thought [73], which strengthens models’ common sense reasoning performance. On top of learning from training data used in pretrained language models, researchers also incorporate external knowledge to improve models’ common sense reasoning ability. For example, Lv et al. [74] leveraged heterogeneous knowledge bases like ConceptNet and Wikipedia and built knowledge graphs to learn graph-based contextual word representation.
Theory of Mind (ToM) [75] is a crucial aspect of reasoning in social contexts. ToM refers to the ability to attribute mental states—beliefs, intentions, desires, emotions—to others and to understand that these mental states may differ from one’s own. This cognitive capacity is fundamental for social interaction and communication. Baker et al. [76] proposed a Bayesian Theory of Mind (BToM) framework that models joint inference of beliefs and desires, demonstrating how people can rationally infer others’ mental states from observed actions. Sap et al. [77] developed ATOMIC, a commonsense knowledge graph that captures social and psychological aspects of everyday events, including people’s intents and reactions. Another work [78] proposed a neural network approach to model people’s mental states from textual narratives. They demonstrated that their model could improve the understanding of social phenomena in stories. These works showcase the potential of integrating Theory of Mind into AI systems for better social understanding.
II-C3 Discussion

### v2

@@ 1 Introduction
Social intelligence of machines to understand social interactions can be the foundation for developing various AI applications. Embodied agents [2] or virtual agents [3] require social understanding capabilities to effectively interact and collaborate with humans. More broadly, capabilities for social understanding can enhance diverse types of AI agents, from chatbots [4] to recommendation systems [5], to better understand user needs and build rapport with humans. By integrating social understanding into systems, we can create more effective and socially adept AI applications across diverse domains.
@@
    Understanding multi-party dynamics: Social interactions often involve multiple participants. Analyzing group conversations requires aggregating individual social behaviors and modeling the interpersonal connections between people. This entails multifaceted aspects of analyzing dynamics such as tracking who is speaking to whom, who is interrupting whom, and who is being positive to whom. Based on the interpretation of such relationships, we can effectively understand social interactions among multiple people.
@@
As a result, the roadmap for ideal social AI involves building machines with capabilities across those three fronts: multimodal understanding, multi-party modeling, and belief awareness. Figure 1 shows the dynamics of social interactions including these capabilities. Through these, machines can approach human social intelligence for seamlessly interpreting social interactions. While progress has been made in narrow applications, advanced social intelligence requires integrating breakthroughs across those areas.
We classify existing works related to understanding social interactions from the perspective of social cue types (i.e., verbal and non-verbal cues). This classification is associated with how to extract social meaning from each social cue, which can provide a foundation for holistically understanding social interactions. Figure 2 presents a taxonomy of existing research on social understanding, organized according to the types of social cues. The taxonomy includes three main branches: verbal cues, non-verbal cues, and multimodal cues. Under the verbal cues branch, key research areas include dialogue act analysis for recognizing intent, dialogue emotion analysis, and commonsense reasoning to interpret socially common knowledge in language. The non-verbal cues branch encompasses research on recognizing non-verbal behaviors including gestures, gaze, and facial expressions. Finally, the multimodal branch covers works that integrate multimodal signals to more comprehensively infer emotions, conversation dynamics, and social situations.
@@
The components listed under each category in Figure 2 are not meant to be isolated research areas. They represent key building blocks and capabilities that need to be connected and integrated to enable truly holistic social understanding. For example, dialogue act analysis and emotion recognition from facial expressions could be combined to better interpret the intent and sentiment behind an utterance. Similarly, gesture recognition and gaze analysis could be jointly leveraged with spoken language to recognize who an utterance is directed at in a multi-party conversation. Therefore, while organizing the existing research, the taxonomy also points towards the need for further integration across areas. In the course of our survey, we will touch upon this need for integration of multimodal social cues, as well as discuss how multi-party dynamics and belief modeling intersect with the social understanding landscape.
@@
This is the first survey that addresses a comprehensive overview of machine learning studies for social understanding, covering both verbal and non-verbal approaches. While there have been surveys addressing specific aspects of social understanding such as gaze analysis [13], dialogue systems [14], and facial expression recognition [15], these works focus on narrow subdomains without considering the broader social context. There also have been position papers on social AI regarding data infrastructure [16] and social AI agents [17]. While these papers provide valuable high-level discussions of relevant works, they do not focus on offering a detailed, comprehensive survey of existing literatures. In contrast, our paper offers a holistic survey of the social understanding landscape, encompassing techniques across multiple modalities and identifying key capabilities crucial for advancing socially intelligent AI. By organizing the wide range of technical works around these core capabilities, we aim to provide a unified perspective and illuminate promising research directions. The scope of our survey covers the problem of social understanding rather than the creation of agents, as we argue that social understanding is a fundamental prerequisite for socially intelligent agents.
@@
We structure our paper as follows. In Sections 2-4, we provide an overview of recent advancements related to understanding verbal, non-verbal, and multimodal social cues, respectively. Section 5 discusses open challenges and promising future research directions based on the suggested key capabilities. Finally, we summarize the key points of this survey and offer concluding remarks in Section 6. By reviewing existing techniques and limitations, we aim to clarify the development trajectory and road ahead for social understanding research.
2 Understanding Verbal Cues
2.1 Dialogue Act Analysis
2.1.1 Social Background
A dialogue act is “a combination of a communicative function and a semantic content” [18]. It refers to the intended action and communicative purpose of an utterance. For instance, when asking for help to open a door, various utterances can fulfill the same dialogue act. This analysis is crucial for social AI systems in natural language understanding, intent detection, and contextual awareness.
@@ 2.1.1 Social Background
2.1.2 Dialogue Act Classification
Early work treated dialogue act classification as a sequence labeling problem and applied artificial neural networks for dialogue act prediction. Lee and Dernoncourt [30] proposed a model based on CNNs and RNNs. Later work used hierarchical neutral networks for the task. More recent work incorporates context-aware self attention mechanism to the hierarchical neural network to further enhance dialogue act understanding. Raheja and Tetreault [31] combined self-attention, hierarchical neural networks and contextual dependencies to develop a new model that leads to performance improvements. Li et al. [32] recognized the association between dialogue acts and subject matters and introduced dual-attention hierarchical RNN to capture such dependency. Chapuis et al. [33] used a hierarchical encoder based on transformer architectures to learn generic representations of dialogues in the pretraining stage. Colombo et al. [34] proposed a sequence-to-sequence model with a hierarchical encoder, guided attention mechanism and beam search method. Moreover, work like [35] considered the interactive nature of dialogue by incorporating turn changes among speakers into dialogue act modeling. Malhotra et al. [29] also introduced speaker- and time-aware contextual learning in their counseling conversation dataset. As such, information about speakers can be captured and it provides more semantic information about the conversations. Furthermore, understanding individuals’ beliefs helps with dialogue state tracking. Mrksic et al. [36] leveraged this idea with a neural belief tracker to predict goals in dialogues.
@@ 2.1.2 Dialogue Act Classification
2.1.3 Discussion
Dialogue act analysis is important for understanding dialogue contexts and the information exchange among different parties. However, it often tends to focus on the local context of the dialogue exchange itself, without fully accounting for the broader social context in which the interaction is embedded. Social factors like power relations and cultural norms can shape dialogue acts in important ways that may be missed. Additionally, much of the research has focused on the verbal aspects of interaction, despite face-to-face conversations involving rich multimodal signals like gaze and gesture. Furthermore, most dialogue act analysis works are based on task-oriented or single-topic interactions between two participants, whereas real-world social interactions often involve multiple participants discussing various topics and goals, adding complexity. Analyzing dialogues with multiple participants further complicates the task as these often lack clear turn-taking and involve overlapping speech, making it difficult to identify dialogue acts well.
@@ 2.1.3 Discussion
2.2 Dialogue Emotion Analysis
2.2.1 Social Background
@@
2.2.2 Dialogue Emotion Recognition
Based on Poria et al. [44], the essential information required to accurately detect the emotion in an utterance from a dialogue can be categorized into three types: 1) the utterance and its surrounding context; 2) the speaker’s state; and 3) emotions in preceding utterances.
Revolving around these three identified categories, researchers have come up with different methods to identify emotions embedded in conversations. Poria et al. applied bidirectional LSTM and attention mechanism to model context from surrounding utterances and Hazarika et al. introduced conversational memory network to exploit speaker memories to model the dialogue context [45]. These early works did not differentiate between various speakers. In light of this, Majumder et al. proposed DialogueRNN [46], which treats each party involved in a dialogue individually and adapts to the speaker of every utterance. Li et al. [47] framed the task under a multi-task setting by introducing an auxiliary task of speaker identification to improve emotion recognition capability in models. Besides individual speaker attributes, modeling inter-speaker relations is also very important. Ghosal et al. [48] introduced DialogueGCN with a graph network to model inter-speaker dependency. On top of inter-speaker relation modeling, researchers also use graph neutral networks to model dialogue context such as local and global context [49], as well as turn modeling [50]. Furthermore, some work incorporated common sense knowledge to enhance emotion recognition in dialogues. Ghosal et al. [51] introduced the COSMIC framework to incorporate common sense knowledge such as causal relations to better understand interactions and emotion tones. Xie et al. [52] introduced Knowledge-Interactive network which utilizes common sense knowledge and sentiment lexison to enrich semantic information. More recently, Lei et al. [53] framed emotion recognition in conversation to a generative plug-and-play framework using Large Language Models, which achieves state-of-the-art performance.
2.2.3 Discussion
@@
2.3 Common Sense Reasoning
2.3.1 Social Background
@@
Common sense reasoning helps to explain the mechanisms of human mind and behavior [56, 57]. It also aids individuals explore the surroundings and understand the social context [58]. During social interactions, common sense reasoning allows individuals to make accurate inferences about others’ beliefs based on observable evidence and social context. Ambiguity in language remains an important challenge for language understanding and common sense reasoning is one key in resolving such ambiguity [59].
2.3.2 Common Sense Reasoning
@@
In the past decade, neural methods have demonstrated remarkable progress in common sense reasoning tasks. In early work, embedding models such as word2vec [60] and GloVe [61] are trained to learn representations of different words and entities in the vector space. Such word representations can help models understand the relationships between different concepts, enhancing their common sense reasoning capabilities. Model architectures like memory networks [62] and Recurrent Entity Network [63] can keep track of the world state with long-term memory. In more recent years, the transformer architecture has been applied in the Generative Pre-trained Transformer (GPT) [64] and the Bidirectional Encoder Representations from Transformers (BERT) [65]. With such contextual language models, the meaning of a word can be different and hence is better captured based on the surrounding context, leading to stronger common sense reasoning capabilities in models. These pretrained models can be further fine-tuned for downstream common sense reasoning tasks to become more task-specific [66]. With increased scale in language models, recent large language models can perform common sense reasoning in zero-shot or few-shot [67] manners. Moreover, Wei et al. introduced step-by-step thinking through Chain of Thought [68], which strengthens models’ common sense reasoning performance. On top of learning from training data used in pretrained language models, researchers also incorporate external knowledge to improve models’ common sense reasoning ability. For example, Lv et al. [69] leveraged heterogeneous knowledge bases like ConceptNet and Wikipedia and built knowledge graphs to learn graph-based contextual word representation.
Theory of Mind (ToM) [70] is a crucial aspect of reasoning in social contexts. ToM refers to the ability to attribute mental states—beliefs, intentions, desires, emotions—to others and to understand that these mental states may differ from one’s own. This cognitive capacity is fundamental for social interaction and communication. Baker et al. [71] proposed a Bayesian Theory of Mind (BToM) framework that models joint inference of beliefs and desires, demonstrating how people can rationally infer others’ mental states from observed actions. Sap et al. [72] developed ATOMIC, a commonsense knowledge graph that captures social and psychological aspects of everyday events, including people’s intents and reactions. Another work [73] proposed a neural network approach to model people’s mental states from textual narratives. They demonstrated that their model could improve the understanding of social phenomena in stories. These works showcase the potential of integrating Theory of Mind into AI for better social understanding.
2.3.3 Discussion